# GP+: A Python Library for Kernel-based learning via Gaussian Processes

### v1

@@ A.1 Engineering Examples
Hybrid organic–inorganic perovskite (HOIP) is a material design problem where the goal is to identify the composition with the smallest inter-molecular binding energy [120]. The dataset used here is generated via three distinct sources that simulate the band gap property of HOIP as a function of composition based on the density functional theory (DFT). These compositions are characterized via three categorical variables that have $10$, $3$, and $16$ levels (i.e., there are $480$ unique compositions in total). The major differences between the datasets generated by the three sources are their fidelity (or accuracy) and size. Specifically, dataset generated by Source $1$ is the most accurate among the three and contains $480$ samples while the datasets generated by Source $2$ and Source $3$ contain $179$ and $240$ samples, respectively, and have lower levels of accuracy compared to Source $1$. In Section 5.1 we exclusively use the HF data but use all three sources with the corresponding costs of $40$, $10$, and $1$ in Section 5.4.
High entropy alloy (HEA) is another alloy design problem with the goal of finding the compound with the lowest thermal expansion coefficient [121]. This is a $5$-dimensional example where the features show the percentage of each element ($Fe,Ni,Co,Cr,V,Cu$) in the alloy. It has one high- and one low-fidelity data sources whose corresponding costs are $50$ and $10$, respectively. Both datasets have $700$ samples and are single-response.
@@
NTA is the problem of designing a nanolaminate ternary alloy that is used in applications such as high-temperature structural materials [122]. These alloys have compositions of the form $M_{2}AX$ where $M$ is an early transition metal, $A$ is a main group element, and $X$ is either carbon or nitrogen. NTA is $3-$dimensional with just categorical features which have $10$, $12$, and $2$ levels. The dataset has $224$ samples and the response is the bulk modulus of the alloy.
@@ A.2 Analytic Examples
To demonstrate the interpretability of embeddings learnt by GP+, we now use Borehole-Mixed  to generate a training dataset of size $400$ and then fit an emulator to it. Upon training, we visualize the latent points learnt for the two categorical variables and color-code them based on either the combination or response magnitude, see Figures 17(a) and 17(b), respectively. As it can be observed in these figures, the learnt embeddings preserve the underlying numerical relations even though GP+ does not have access to the numerical values used in data generation. For instance, all the variability in the latent space is in two directions with is in line with the fact that all combinations of the categorical data can be quantified via the two underlying numerical features. We also observe that close-by (distant) latent points have similar (different) response values which increases as we move from the top-right corner to the bottom left corner of the latent space, see Figure 17(b). These embeddings also indicate variable importance. For instance, based on Figure 17(a) we observe that changing the levels of the first categorical variable from "a" to "e" results in larger latent movements than changing the levels of the second categorical variable. Such a behavior indicates the the underlying function is more sensitive to the former variable and we validate this argument in Appendix B by conducting sensitivity analyses.
@@ A.3 Calibration Examples
Wing, Borehole and beam deflection are used in Section 5.3 to illustrate the performance of GP+ in calibration. Their formulation and details are provided in Table 5. As explained in Section 5.3, there are $2$ and $4$ calibration parameters in Borehole and Wing, respectively. Also, beam deflection is a bi-fidelity $5$-dimensional problem [106] whose features are a constant distributed load applied to the beam ($p=12000N/m$), width ($b=0.15m$), height ($h=0.3m$), and the length of the beam ($L=5\leavevmode\nobreak\ m$). The last feature is the unknown material Young’s modulus ($\zeta$) which we aim to estimate. The analytic formulation and more details regarding beam deflection is also presented in Table 5.
@@ Appendix B Sensitivity Analysis
Sobol sensitivity analysis is a global variance-based method used for quantifying each input’s main and total contribution to the output variance [123]. While main-order Sobol indices (SIs) reveal the individual contributions of input variables, total-order indices capture both the individual and interaction effects of inputs on the output. We highlight that while Sobol indices are typically applied to quantitative features, we extend the idea to qualitative features in GP+ by sampling random quantitative values and mapping them to the unique levels of a categorical variable. This functionality is accessible with the model.Sobol() command in GP+.
@@ Appendix C Emulation and Optimization Options
U_train Input Tensor
y_train Output (responce) Tensor
qual_index Column index and number of levels of categorical variables {}
@@
lv_dim Dimension of manifold to be learnt $2$
@@
fix_noise Flag for estimating noise via the nugget parameter False
noise Fixed noise value if fix_noise = True $41e-$ 4
lb_noise Lower bound for noise $81e-$ 8
@@
encoding_type Type of $f_{\boldsymbol{\pi}}(\boldsymbol{t})$one-hot"
manifold_type Type of embedding "deterministic"
uniform_encoding_columns Number of columns for uniform encoding $2$
lv_columns Which categorical features be learned separately [ ]
basis Type of mean function for GP "single"
NN_layers_mean Structure of neural network for mean function $[4,4]$
@@
Table 9: Model initialization options: We provide a range of options that streamline the adoption of GP+ in many applications with just a few lines of codes.
@@ Appendix E Single-Task Multi-Fidelity GP (STMF-GP)
The STMF-GP modifies the correlation function of GP for MF hierarchical MF modeling. Specifically, it adopts an additive covariance function that relies on introducing two user-defined quantitative features [6, 30]. The first feature, denoted by $x_{a}$, is restricted to the $[0,1]$ range and assigns a fidelity value to a source based on the user’s belief (larger values correspond to higher fidelities). This assigned fidelity value directly affects the correlation and cost function. The second feature, denoted by $x_{b}$, is the iteration fidelity parameter and benefits MF BO specifically in the context of hyperparameter tuning of large machine learning models. These two features are used in three user-defined functions defined as follows. $e_{1}(\cdot)$ and $e_{3}(\cdot)$ are bias kernels that aim to take the discrepancies among the sources into account:
@@
Finally, the modified covariance function is [124]:
@@ G.1 BO Improvements
As mentioned in Section 5.4, the accuracy of the emulator in quantifying uncertainties significantly affects the performance of BO. This impact is more noticeable in MF problems where biased LF sources can misguide the optimization process. To address this challenge, we employ interval scores ($IS$) to penalize the objective function. We choose $IS$ since it is robust to outliers, rewards narrow prediction intervals, and is flexible in the choice of desired coverage levels [125, 126]. $IS$ is a special case of quantile prediction that penalizes the emulator for each observation that is not inside the $(1-v)\times 100\%$ prediction interval and is calculated as:
@@
where ||$\cdot$|| ⋅ | denotes the absolute function and $\epsilon$ is a user-defined scaling parameter which is set to $0.08$ by default in GP+. We refer the reader to [87] for more in-depth information.
@@ G.2 BO Options
Table 13: Options provided by GP+ for Bayesian optimization: GP+ accommodates both single- and multi-fidelity BO. The user can provide both analytic functions and dataset where the former is typically used in comparison studies while the latter is used in real world applications.
@@ Appendix H BoTorch
BoTorch is an MF cost-aware BO package that employs STMF-GP (explained in Appendix E) as the emulator and leverages the knowledge gradient (KG) as the AF. KG is a look-ahead AF that chooses the next sampling point ($\boldsymbol{x}^{(k+1)}$) based on the effect of the yet-to-be-seen observation (i.e., $y^{k+1}$ which follows a normal distribution) on the optimum value predicted by the emulator. This AF quantifies the expected utility of $\boldsymbol{x}$ at iteration $k+1$ as:
@@
where $\mathcal{D}^{k}=\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{k}$ is the training data in iteration $k$ and $\max{\mu^{(k)}}=\max{\mu^{(k)}}(\boldsymbol{x})$ denotes the maximum mean prediction of the emulator trained on $\mathcal{D}^{k}$. The expectation operation in Equation H-10 appears due to the fact that $y^{(k+1)}$ is not observed yet and $\alpha_{KG}(\boldsymbol{x})$ is relying on the predictive distribution provided by the emulator that is trained on $\mathcal{D}^{k}$. This expectation cannot be calculated analytically and hence a Monte Carlo estimate is used in practice:
@@
where $\max{{\mu^{(k+1)}}^{m}}=\max{{\mu^{(k+1)}}^{m}(\boldsymbol{x})}$ is calculated by first drawing a sample at $\boldsymbol{x}$ from the emulator that is trained on $\mathcal{D}^{k}$ and then retraining the emulator on $\mathcal{D}^{k}\cup(\boldsymbol{x},y^{m})$ where $y^{m}$ is response of the drawn sample. In practice, a small value must be chosen for $M$ since maximizing $\gamma_{KG}(\boldsymbol{x})$ over the input space at each iteration of BO is very expensive. Refer to [127, 128] for more information on KG and its implementation.

### v2

@@ A.1 Engineering Examples
Hybrid organic–inorganic perovskite (HOIP) is a material design problem where the goal is to identify the composition with the smallest inter-molecular binding energy [egger2016hybrid]. The dataset used here is generated via three distinct sources that simulate the band gap property of HOIP as a function of composition based on the density functional theory (DFT). These compositions are characterized via three categorical variables that have $10$, $3$, and $16$ levels (i.e., there are $480$ unique compositions in total). The major differences between the datasets generated by the three sources are their fidelity (or accuracy) and size. Specifically, dataset generated by Source $1$ is the most accurate among the three and contains $480$ samples while the datasets generated by Source $2$ and Source $3$ contain $179$ and $240$ samples, respectively, and have lower levels of accuracy compablack to Source $1$. In Section 5.1 we exclusively use the HF data but use all three sources with the corresponding costs of $40$, $10$, and $1$ in Section 5.4.
High entropy alloy (HEA) is another alloy design problem with the goal of finding the compound with the lowest thermal expansion coefficient [rao2022machine]. This is a $5$-dimensional example where the features show the percentage of each element ($Fe,Ni,Co,Cr,V,Cu$) in the alloy. It has one high- and one low-fidelity data sources whose corresponding costs are $50$ and $10$, respectively. Both datasets have $700$ samples and are single-response.
@@
NTA is the problem of designing a nanolaminate ternary alloy that is used in applications such as high-temperature structural materials [cover2009comprehensive]. These alloys have compositions of the form $M_{2}AX$ where $M$ is an early transition metal, $A$ is a main group element, and $X$ is either carbon or nitrogen. NTA is $3-$dimensional with just categorical features which have $10$, $12$, and $2$ levels. The dataset has $224$ samples and the response is the bulk modulus of the alloy.
@@ A.2 Analytic Examples
To demonstrate the interpretability of embeddings learnt by GP+, we now use Borehole-Mixed  to generate a training dataset of size $400$ and then fit an emulator to it. Upon training, we visualize the latent points learnt for the two categorical variables and color-code them based on either the combination or response magnitude, see Figures 19(a) and 19(b), respectively. As it can be observed in these figures, the learnt embeddings preserve the underlying numerical relations even though GP+ does not have access to the numerical values used in data generation. For instance, all the variability in the latent space is in two directions with is in line with the fact that all combinations of the categorical data can be quantified via the two underlying numerical features. We also observe that close-by (distant) latent points have similar (different) response values which increases as we move from the top-right corner to the bottom left corner of the latent space, see Figure 19(b). These embeddings also indicate variable importance. For instance, based on Figure 19(a) we observe that changing the levels of the first categorical variable from "a" to "e" results in larger latent movements than changing the levels of the second categorical variable. Such a behavior indicates the the underlying function is more sensitive to the former variable and we validate this argument in Appendix B by conducting sensitivity analyses.
@@ A.3 Calibration Examples
Wing, Borehole and beam deflection are used in Section 5.3 to illustrate the performance of GP+ in calibration. Their formulation and details are provided in Table 5. As explained in Section 5.3, there are $2$ and $4$ calibration parameters in Borehole and Wing, respectively. Also, beam deflection is a bi-fidelity $5$-dimensional problem [marelli2014uqlab] whose features are a constant distributed load applied to the beam ($p=12000N/m$), width ($b=0.15m$), height ($h=0.3m$), and the length of the beam ($L=5\leavevmode\nobreak\ m$). The last feature is the unknown material Young’s modulus ($\zeta$) which we aim to estimate. The analytic formulation and more details regarding beam deflection is also presented in Table 5.
@@ Appendix B Sensitivity Analysis
Sobol sensitivity analysis is a global variance-based method used for quantifying each input’s main and total contribution to the output variance [saltelli2010variance]. While main-order Sobol indices (SIs) reveal the individual contributions of input variables, total-order indices capture both the individual and interaction effects of inputs on the output. We highlight that while Sobol indices are typically applied to quantitative features, we extend the idea to qualitative features in GP+ by sampling random quantitative values and mapping them to the unique levels of a categorical variable. This functionality is accessible with the model.Sobol() command in GP+.
@@ Appendix C Emulation and Optimization Options
train_x Input Tensor
train_y Output (response) Tensor
dtype Data type of the model and data torch.float
device Device to build the model (CPU or CUDA) "cpu"
qual_dict Column index and number of levels of categorical variables {}
@@
lb_noise Lower bound for noise $81e-$ 8
fix_noise Flag for estimating noise via the nugget parameter False
fix_noise_val Fixed noise value if fix_noise = True $51e-$ 5
@@
fixed_length_scale Flag to fix the length scale False
fixed_length_scale_val Fixed length scale value torch.tensor([1.0])
encoding_type Type of $f_{\boldsymbol{\pi}}(\boldsymbol{t})$one-hot"
embedding_dim Dimension of embedding (manifold) to be learnt $2$
separate_embedding Which categorical features be learned separately [ ]
embedding_type Type of embedding "deterministic"
@@
m_gp Type of mean function for GP "single"
m_gp_ref Mean function for reference source (ID = 0) "zero"
NN_layers_m_gp Structure of neural network for mean function $[4,4]$
@@
mean_prior_cal Mean prior for calibration parameter 00
std_prior_cal Standard deviation prior for calibration parameter $1$
interval_score Interval scoring during optimization False
num_pass_train Number of training passes; deterministic/probabilistic $1/20$
num_pass_pblack Number of pblackiction passes; deterministic/probabilistic $1/30$
Table 9: Model options: We provide a range of options that streamline the adoption of GP+ in many applications with just a few lines of code.
@@ Appendix E Single-Task Multi-Fidelity GP (STMF-GP)
The STMF-GP modifies the correlation function of GP for MF hierarchical MF modeling. Specifically, it adopts an additive covariance function that relies on introducing two user-defined quantitative features [RN1392, RN1270]. The first feature, denoted by $x_{a}$, is restricted to the $[0,1]$ range and assigns a fidelity value to a source based on the user’s belief (larger values correspond to higher fidelities). This assigned fidelity value directly affects the correlation and cost function. The second feature, denoted by $x_{b}$, is the iteration fidelity parameter and benefits MF BO specifically in the context of hyperparameter tuning of large machine learning models. These two features are used in three user-defined functions defined as follows. $e_{1}(\cdot)$ and $e_{3}(\cdot)$ are bias kernels that aim to take the discrepancies among the sources into account:
@@
Finally, the modified covariance function is [RN1790]:
@@ G.1 BO Improvements
As mentioned in Section 5.4, the accuracy of the emulator in quantifying uncertainties significantly affects the performance of BO. This impact is more noticeable in MF problems where biased LF sources can misguide the optimization process. To address this challenge, we employ interval scores ($IS$) to penalize the objective function. We choose $IS$ since it is robust to outliers, rewards narrow pblackiction intervals, and is flexible in the choice of desiblack coverage levels [bracher2021evaluating, mitchell2017proper]. $IS$ is a special case of quantile pblackiction that penalizes the emulator for each observation that is not inside the $(1-v)\times 100\%$ pblackiction interval and is calculated as:
@@
where ||$\cdot$|| ⋅ | denotes the absolute function and $\epsilon$ is a user-defined scaling parameter which is set to $0.08$ by default in GP+. We refer the reader to [foumani2023effects] for more in-depth information.
@@ G.2 BO Options
Table 13: Options provided by GP+ for Bayesian optimization: GP+ accommodates both single- and multi-fidelity BO. The user can provide both analytic functions and datasets where the former is typically used in comparison studies while the latter is used in real world applications.
@@ Appendix H BoTorch
BoTorch is an MF cost-aware BO package that employs STMF-GP (explained in Appendix E) as the emulator and leverages the knowledge gradient (KG) as the AF. KG is a look-ahead AF that chooses the next sampling point ($\boldsymbol{x}^{(k+1)}$) based on the effect of the yet-to-be-seen observation (i.e., $y^{k+1}$ which follows a normal distribution) on the optimum value pblackicted by the emulator. This AF quantifies the expected utility of $\boldsymbol{x}$ at iteration $k+1$ as:
@@
where $\mathcal{D}^{k}=\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{k}$ is the training data in iteration $k$ and $\max{\mu^{(k)}}=\max{\mu^{(k)}}(\boldsymbol{x})$ denotes the maximum mean pblackiction of the emulator trained on $\mathcal{D}^{k}$. The expectation operation in Equation H-10 appears due to the fact that $y^{(k+1)}$ is not observed yet and $\alpha_{KG}(\boldsymbol{x})$ is relying on the pblackictive distribution provided by the emulator that is trained on $\mathcal{D}^{k}$. This expectation cannot be calculated analytically and hence a Monte Carlo estimate is used in practice:
@@
where $\max{{\mu^{(k+1)}}^{m}}=\max{{\mu^{(k+1)}}^{m}(\boldsymbol{x})}$ is calculated by first drawing a sample at $\boldsymbol{x}$ from the emulator that is trained on $\mathcal{D}^{k}$ and then retraining the emulator on $\mathcal{D}^{k}\cup(\boldsymbol{x},y^{m})$ where $y^{m}$ is response of the drawn sample. In practice, a small value must be chosen for $M$ since maximizing $\gamma_{KG}(\boldsymbol{x})$ over the input space at each iteration of BO is very expensive. Refer to [frazier2008knowledge, balandat2020botorch] for more information on KG and its implementation.
Appendix I Computational Costs
[…]
In numerical examples (Wing and Borehole, see Figure 21), the formulations of all methods are the same and they only differ in the parameter estimation and optimization. Specifically, for MATLAB, we employ default settings that utilize MLE for parameter estimation and the BFGS algorithm for optimization. The lower computational costs observed in MATLAB are attributed to its optimization process and streamlined implementation. SMT2 exhibits slightly lower computational costs compablack to GP+ and GPytorch which is due to its use of the profiling method. Furthermore, while GP+ and GPytorch are quite similar, the minor difference in their computational costs stems from the different optimizers they employ (GP+ uses L-BFGS from SciPy by default, while GPytorch uses Adam [^20] Adaptive Moment Estimation from PyTorch).
Figure 20: Training cost in examples with categorical inputs: Each bar in the graph represents the computational costs associated with different methods for various examples. Since EHH, HH, and Gower are from SMT2, we use the same color to represent them. In the context of mixed-input problems, the manifold learning process of GP+ leads to a slight increase in its computational cost. However, given the significantly higher accuracy it achieves, this difference in time is justified.
Figure 21: Training cost in examples with all-numerical inputs: Each bar in the graph represents the computational costs associated with different methods for various examples. In the context of numerical problems, all the benchmarks have the same formulation and they just differ in their optimization. This difference causes slightly different computational costs.
[…]